estion 10 What is the pH of 5.69 * 10-2 M HNO 3(aq)? a. 2.54 b.5.65 c. 1.80 O d. 1.24 e.-1.01
Answer #1

- MNOG is strong acid . HNO - 5.69x16-2 HT + Nos 5.69 x 10-2 we know that pH = -log [hty pH = -log (5.698102) .PL = -109(5.69